OpenStudy (anonymous):

Someone help me with the 2 problems I posted yesterday and didn't get answers to... (they're on my profile)

OpenStudy (anonymous):

Ykz can you help or no?

OpenStudy (anonymous):

I get the idea of plugging in one for the other but the problems look weird because they are fractional.

OpenStudy (anonymous):

phi can u help?

OpenStudy (phi):

\[f \circ g= f(g(x))\]\[g(x)= \frac{-1-5x}{5x}\]\[f(x)=\frac{-1}{5x+5}\]

OpenStudy (anonymous):

so far so good

OpenStudy (anonymous):

get on the chat thing Foster

OpenStudy (phi):

The easiest way to do this is concentrate on just the denominator 5x+5 in f(x) replace the x with g(x): \[5\cdot \frac{-1-5x}{5x}+5 \] simplify by canceling the 5's. Also, the make a common denominator for the second 5: \[\frac{-1-5x}{x}+\frac{5}{1}\cdot \frac{x}{x}= \frac{-1-5x+5x}{x}= \frac{-1}{x}\]

OpenStudy (phi):

Now we have the denominator of f(g(x)) simplified to -1/x. So the total expression is \[\frac{-1}{\frac{-1}{x}}= -1\cdot \frac{x}{-1}=x\]

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

shouldnt f(g(x)) = f(-1 / 5(-1-5x/5x)+5)

OpenStudy (anonymous):

when u plug in one for the other

OpenStudy (phi):

what f(g(x)) means is look at the f function, and everywhere you see x, replace it with g(x) so, without expanding it, we would get: \[f(g(x))= \frac{-1}{5g(x)+5}\] of course g(x) is a messy fraction, but you get the idea?

OpenStudy (anonymous):

yeah i know that i have to plug in the fraction for g(x)... just need help simplifying

OpenStudy (anonymous):

so that i get x from there

OpenStudy (phi):

OK, first simplify 5g(x)

OpenStudy (anonymous):

do i multiply 5 on the bottom and top of the fraction?

OpenStudy (anonymous):

so i would get -5 -25x / 25x

OpenStudy (phi):

\[5g(x)= 5\cdot \frac{-1-5x}{5x}\]

OpenStudy (phi):

So the 5 in top and the bottom cancel each other out (5/5 = 1)

OpenStudy (anonymous):

so i dont even have to multiply cause they cancel out?

OpenStudy (phi):

If you are following, then you get \[\frac{-1-5x}{x} \]

OpenStudy (phi):

Now add 5.

OpenStudy (anonymous):

foster: delete ur comments and it wont show up

OpenStudy (phi):

It is always good to cancel things. You could multiply things out but it's messier

OpenStudy (anonymous):

yeah i agree

OpenStudy (anonymous):

let me do it on paper and ill get back to you in a sec or 2

OpenStudy (anonymous):

ok and then?

OpenStudy (phi):

You asked above: so i would get -5 -25x / 25x no. when you multiply fractions, you multiply top times top, bottom times bottom In the case of a whole number like 5, you assume the bottom is 1 (5/1 = 5). So, if you did multiply things out it would be \[\frac{5}{1}\cdot\frac{-1-5x}{5x}= \frac{-5-25x}{5x}\]

OpenStudy (anonymous):

yeah i get that

OpenStudy (phi):

Did you add 5 to \[\frac{-1-5x}{x} \] ?

OpenStudy (anonymous):

too many rules to remember sometimes

OpenStudy (phi):

I agree about the rules. lots of them to remember.

OpenStudy (anonymous):

yes and got -1 / 4-5x/x

OpenStudy (phi):

You got \[ \frac{-1}{4} - \frac{5x}{x} \]

OpenStudy (phi):

How?!

OpenStudy (anonymous):

i added 5 to -1 so i got 4

OpenStudy (anonymous):

why? what am i doing wrong?

OpenStudy (phi):

Meanwhile, to add fractions, you make sure both fractions have the same bottom, and then you can add the tops. 5 is a whole number, so its bottom is understood to be 1

OpenStudy (phi):

The first fraction has an x in the bottom. So you have to change the 5 to have an x in its bottom.

OpenStudy (anonymous):

o

OpenStudy (anonymous):

so the same fractional rules apply with variables as well

OpenStudy (phi):

The upside to the rules, no exceptions.

OpenStudy (anonymous):

i see

OpenStudy (anonymous):

i got x :D

OpenStudy (phi):

when you add 5, you change it to 5x/x and now you add the tops (-1-5x+5x)/x = -1/x so you should get -1/x